karleighb23:

A parachutist's rate during a free fall reaches 231 feet per second. What is this rate in meters per second? At this rate, how many meters will the parachutist fall during 15 seconds of free fall? In your computations, assume that 1 meter is equal to 3.3 feet. Do not round your answers.

justjm:

The most coherent way to solve this problem is using dimensional analysis 1. What is this rate in meters per second? \(\frac{231\ \text{ft}}{1\ \sec}\cdot\frac{1\ \text{m}}{3.3\ \text{ft}}=\text{__}\frac{m}{\sec}\) Notice how the feet cancels, to form m/s. You can fill in the blank. 2. At this rate, how many meters will the parachutist fall during 15 seconds of free fall? So, you use your rate you find from question 1, and multiply it with 15 seconds \(\text{__}\frac{\text{m}}{\sec}\cdot15\ \sec\) The most important part in forming these equations is noticing how the units cancel and get replaced with the goal unit. Got that @karleighb23?

karleighb23:

still confused a little @justjm

justjm:

Okay, let's stick to the first question \(\frac{231\ \text{ft}}{1\ \sec}\cdot\frac{1\ \text{m}}{3.3\ \text{ft}}=\frac{231\ \text{ft}\ \cdot\ 1\ \text{m}}{1\ \sec\ \cdot\ 3.3~\text{ft}}\) Understood that? Now, since we have like units in the numerator and denominator, we can cancel them. When I say cancel, what I really mean to say is that \(\frac{\text{ft}}{\text{ft}}=1\) By this, we have ft. on the numerator and denominator, so those units can be canceled, correct? \(\frac{231\ m}{1\ \sec\ \cdot\ 3.3}\) Now you can simplify That's for question 1

justjm:

Are you there @karleighb23?